# SAGMAN: Stability Analysis of Graph Neural Networks on the Manifolds

### v3

@@ 4.5 SAGMAN-guided GNN Stability Enhancement
To improve the stability of a GNN model, a naive approach is to replace the entire input graph with the low-dimensional graph-based manifold for GNN predictions. However, due to the significantly increased densities in the graph-based manifold, the GNN prediction accuracy may also be influenced.
To achieve a flexible tradeoff between model stability and prediction accuracy, we only modify the top few most unstable nodes identified by SAGMAN. Specifically, for the most unstable nodes we only select the inter-cluster edges, as shown in Figure 2 (d), from the input graph-based manifold and insert them into the original graph. Since the inter-cluster (bridge-like) edges are mostly spectrally critical (with high sampling probabilities defined in Equation 8), adding them to the original graph will mostly significantly alter the structural (spectral) properties of the original graph.
Table 4 presents the error rates for both the original and the enhanced graphs. Nettack was employed to perturb SAGMAN-selected samples within each graph. Notably, reintegrating selected edges into the original graph significantly diminished the error rate when subjected to adversarial evasion attacks.
4.6 Runtime Scalability of SAGMAN
Figure 5: SAGMAN’s runtime across datasets
Figure 5 demonstrates the high efficiency of SAGMAN in handling large graph datasets, attributed to its near-linear time complexity.
@@ 5 Conclusions
In this work, we present SAGMAN, a novel framework for analyzing the stability of GNNs through individual nodes, which is achieved by assessing the resistance-distance distortions incurred by the maps between low-dimensional input and output graph-based manifolds. A key component in SAGMAN is a novel graph dimensionality reduction (GDR) approach for constructing resistance-preserving input/output graph-based manifolds. SAGMAN achieves promising performance in GNN stability analysis, leading to significantly enhanced targeted attacks and GNN stability. The current SAGMAN framework is most effective for handling graphs that can be well represented in low-dimensional space and have large eigengaps.
Appendix A Metrics for Assessing GNN Stability
@@
Appendix B Algorithm Flow of SAGMAN
The Algorithm 1 shows the key steps in SAGMAN. The process of obtaining low-dimensional graphs $G_{X}$ and $G_{Y}$ involves several steps, starting with the reduction of the graph dimension of $G_{X}$ using the weighted spectral embedding matrix. This matrix is denoted as $U_{k}$, which is defined in Definition 3.1. Following the approach suggested by (Deng et al., 2022), we enhance the graph construction by concatenating node features with dominant eigenvectors. The next step involves constructing the input PGM. This is achieved by first constructing a dense graph using the embedding matrix $E$, and then sparsifying this dense graph through short-cycle decomposition. The structure of the low-dimensional graph $G_{X}$ is obtained with the assistance of the input PGM. This graph, along with the node feature $X$, is then fed into a GNN to generate the output matrix $Y$. To construct the output PGM, we follow a similar process as with the input PGM: we construct a dense graph with $Y$ and sparsify the dense graph via short-cycle decomposition. This results in the low-dimensional graph $G_{Y}$, which is obtained with the aid of the output PGM. From the graphs $G_{X}$ and $G_{Y}$, we derive the Laplacian matrices $L_{X}$ and $L_{Y}$, respectively. The metric $\delta^{M}(p,q_{i})$ can then be calculated by evaluating $V_{N}^{\top}e_{p,q_{i}}$. In Appendix F, we provide a more detailed explanation of the relationship between $V_{N}^{\top}e_{p,q_{i}}$ and $\delta^{M}(p,q_{i})$. Finally, the node $p$ DMD score is obtained by averaging the DMD scores of node $p$ and its neighbors.
Algorithm 1 The algorithm flow of SAGMAN
  Input: Input graph $G$, Node features $X$, GNN
  Output: Stability (DMD) scores for all nodes
  $U_{N}\leftarrow\text{compute\_weighted\_spectral\_embedding}(G)$
  $V\leftarrow\text{concatenate}(X,U_{N})$
  $kNN\_dense\_graph\leftarrow\text{construct\_kNN\_graph}(E)$
  $input\_PGM\leftarrow\text{sparse\_graph}(kNN\_dense\_graph)$
  $G_{X}\leftarrow\text{get\_low\_dimensional\_graph}(input\_PGM)$
  $Y\leftarrow\text{GNN}(G_{X},X)$
  $kNN\_dense\_graph\_Y\leftarrow\text{construct\_kNN\_graph}(Y)$
  $output\_PGM\leftarrow\text{sparse\_graph}(kNN\_dense\_graph\_Y)$
  $G_{Y}\leftarrow\text{get\_low\_dimensional\_graph}(output\_PGM)$
  $L_{X}\leftarrow\text{compute\_laplacian}(G_{X})$
  $L_{Y}\leftarrow\text{compute\_laplacian}(G_{Y})$
  $V_{k}\leftarrow\text{calculate\_generalized\_eigenvectors}(L_{X},L_{Y})$
  for each node $p$ in $G_{X}$ do
     Compute the node stability score for $p$ in $G_{X}$:
$score(p)=\frac{1}{|\mathbb{N}_{X}(p)|}\sum_{q_{i}\in\mathbb{N}_{X}(p)}\left(\|V_{k}^{\top}e_{p,q}\|^{2}_{2}\right)$
  end for
Appendix C SAGMAN without GDR

### v4

@@ 4.5 SAGMAN-guided GNN Stability Enhancement
A naive approach to improving the stability of a GNN model is to replace the entire input graph with the low-dimensional graph-based manifold for GNN predictions. However, due to the significantly increased densities in the graph-based manifold, the GNN prediction accuracy may be adversely affected. To achieve a flexible trade-off between model stability and prediction accuracy, we select only the inter-cluster edges from the input graph-based manifold, as shown in Figure 2(d), and insert them into the original graph. Since the inter-cluster (bridge-like) edges are typically spectrally critical—with high sampling probabilities as defined in Theorem 3.2—adding them to the original graph significantly alters its structural (spectral) properties.
Whereas previous work (Deng et al., 2022) focuses on improving the robustness of GNNs against poisoning attacks, our work centers on enhancing robustness against evasion attacks. In Table 4, we present the error rates for both the original and the enhanced graphs, where Nettack was employed to perturb SAGMAN-selected most unstable samples within each graph. Table 5 shows the error rates for the state-of-the-art GOOD-AT (Li et al., 2024) method and our proposed SAGMAN approach. Since SAGMAN is a versatile plug-in method that can be combined with other robustness techniques, we also report results for the combined application of GOOD-AT and SAGMAN. Notably, reintegrating selected edges into the original graph significantly reduces the error rate when subjected to adversarial evasion attacks.
@@ 5 Conclusions
In this work, we introduced SAGMAN, a novel framework for analyzing the stability of GNNs at the individual node level by assessing the resistance-distance distortions between low-dimensional input and output graph-based manifolds. A key component of SAGMAN is the proposed Graph Dimensionality Reduction (GDR) approach for constructing resistance-preserving manifolds, which enables effective stability analysis.
Our experimental results demonstrate that SAGMAN effectively quantifies GNN stability, leading to significantly enhanced targeted adversarial attacks and improved GNN robustness. The current SAGMAN framework is particularly effective for graphs that can be well represented in low-dimensional spaces and exhibit large eigengaps.
Appendix A Appendix
A.1 Graph-based Manifolds and Distance Metrics
Graph-based Manifolds.
A manifold is a topological space that locally resembles Euclidean space near each point (Lee, 2012). In our work, we utilize graph-based manifolds to represent complex data structures. Specifically, we represent each manifold as an undirected (connected) graph $G=(V,E)$, where $V$ is the set of vertices corresponding to data points, and $E$ is the set of edges encoding relationships (e.g., conditional dependencies) between these points (Tenenbaum et al., 2000). This representation is particularly effective when the underlying data structure can be approximated by a network of discrete points, as is common in spectral clustering and manifold learning (Belkin & Niyogi, 2003).
Mappings Between Manifolds.
Understanding the mappings between manifolds is essential for transferring and comparing information across different data representations. In our context, these mappings correspond to transformations between the input and output graph-based manifolds of an ML model. Formally, a mapping $\varphi:M_{X}\rightarrow M_{Y}$ from the input manifold $M_{X}$ to the output manifold $M_{Y}$ allows us to analyze how the model transforms data points and their relationships (Pan & Yang, 2009). This is crucial for assessing the model’s stability with respect to small perturbations in the input.
Distance Metrics on Manifolds.
Measuring dissimilarities between points on manifolds requires appropriate distance metrics. While geodesic distances, representing the shortest paths between points on a manifold, are commonly used, they can be computationally intensive for large datasets and may not capture global structural information due to their inherently local nature (Tenenbaum et al., 2000).
Effective Resistance Distances on Graph-based Manifolds.
To overcome these limitations, we employ the effective resistance distance from electrical network theory. By modeling the undirected graph as a resistive network, the effective resistance between two nodes captures both local relationships and global structural properties of the graph (see Appendix A.1 for more details). This metric effectively combines the advantages of geodesic and global distances, providing a more comprehensive measure of similarity in graph-based manifolds (Tenenbaum et al., 2000). Additionally, it is computationally efficient for large-scale graphs (Klein & Zhu, 1998), making it practical for our analysis.
To illustrate the difference between geodesic distance and effective resistance distance, consider the example in Figure 4. In all three graphs, nodes $A$ and $B$ have the same geodesic distance (i.e., the shortest path length is identical). However, their effective resistance distances $\Omega_{AB}$ vary significantly due to differences in the global structure of each graph. This example demonstrates that while geodesic distance only accounts for the shortest path, effective resistance distance incorporates the overall connectivity and network topology, making it a more informative metric for assessing similarities in graph-based manifolds.
Figure 4: Examples of three different graph structures with nodes $A$ and $B$. Despite having the same geodesic distance (shortest path length), the effective resistance distances $\Omega_{AB}$ vary due to the different global structures of the graphs.
By incorporating effective resistance distances into our stability analysis, we can more accurately assess the distance distortions introduced by the mappings between the input and output manifolds (as discussed in Section 2.1). This provides a solid foundation for evaluating the stability of ML models on graph-based manifolds, enhancing both the theoretical rigor and practical applicability of our approach.
A.2 Spectral Graph Theory
${A}(i,j)=\begin{cases}w(i,j)&\text{ if }(i,j)\in E\\ 0&\text{otherwise }\end{cases}\vspace{-0pt}$ (6)
Lemma A.1.
Lemma A.1 is the Courant-Fischer Minimax Theorem (Golub & Van Loan, 2013) for solving the eigenvalue problem: $Lu_{k}=\lambda_{k}u_{k}$. The generalized Courant-Fischer Minimax Theorem for solving generalized eigenvalue problem $L_{X}v_{k}=\lambda_{k}L_{Y}v_{k}$ can be expressed as follows:
Lemma A.2.
(The Generalized Courant-Fischer Minimax Theorem) Given two Laplacian matrices $L_{X},L_{Y}\in\mathbb{R}^{|V|\times|V|}$ such that $null(L_{Y})\subseteq null(L_{X})$, $L_{Y}^{+}$ denotes the Moore–Penrose pseudoinverse of $L_{Y}$, the $k$-th largest eigenvalue of $L_{Y}^{+}L_{X}$ can be computed under the condition of $1\leq k\leq rank(L_{Y})$ by:
A.3 Graph-based Manifold Learning via PGMs
Given $M$ samples of $N$-dimensional vectors stored in a data matrix $X\in{\mathbb{R}^{N\times M}}$, the recent graph topology learning methods (Kalofolias & Perraudin, 2019; Dong et al., 2019) estimate graph Laplacians from $X$ for achieving the following desired characteristics:
Smoothness of Graph Signals. The graph signals corresponding to the real-world data should be sufficiently smooth on the learned graph structure: the signal values will only change gradually across connected neighboring nodes. The smoothness of a signal $x$ over an undirected graph $G=(V,E,w)$ can be measured with the following Laplacian quadratic form: ${x^{\top}}Lx=\sum\limits_{\left({p,q}\right)\in E}{{w_{p,q}}{{\left({x\left(p\right)-x\left(q\right)}\right)}^{2}}}$, where $w_{p,q}$ denotes the weight of edge ($p,q$), $L=D-W$ denotes the Laplacian, $D$ denotes the diagonal (degree) matrix, and $W$ denotes the adjacency matrix of $G$, respectively. The smaller quadratic form implies the smoother signals across the edges in the graph. The smoothness ($Q$) of a set of signals $X$ over graph $G$ is computed using the following matrix trace (Dong et al., 2019): $Q(X,L)=Tr({X^{\top}LX})$, where $Tr(\bullet)$ denotes the matrix trace.
Sparsity of the Estimated Graph. Graph sparsity is another critical consideration in graph learning. One of the most important motivations of learning a graph is to use it for downstream computing tasks. Therefore, more desired graph topology learning algorithms should allow better capturing and understanding the global structure (manifold) of the data set, while producing sufficiently sparse graphs that can be easily stored and efficiently manipulated in the downstream algorithms, such as circuit simulations, network partitioning, dimensionality reduction, data visualization, etc.
Problem Formulation. Consider a random vector $x\sim N(0,\Sigma)$ with probability density function:
$f(x)=\frac{\exp{\left(-\frac{1}{2}x^{\top}\Sigma^{-1}x\right)}}{(2\pi)^{N/2}\det(\Sigma)^{(1/2)}}\propto\det(\Theta)^{1/2}\exp{\left(-\frac{1}{2}x^{\top}\Theta x\right)},$ (9)
where $\Sigma=\mathbb{E}[xx^{\top}]\succ 0$ denotes the covariance matrix, and $\Theta=\Sigma^{-1}$ denotes the precision matrix (inverse covariance matrix). Prior graph topology learning methods aim at estimating sparse precision matrix $\Theta$ from potentially high-dimensional input data, which fall into the following two categories:
(A) The graphical Lasso method aims at estimating a sparse precision matrix $\Theta$ using convex optimization to maximize the log-likelihood of $f(x)$ (Friedman et al., 2008):
$\max_{\Theta}:\log\det(\Theta)-Tr(\Theta S)-{\beta}{{\|\Theta\|}}_{1},$ (10)
[…]
(B) The Laplacian estimation methods have been recently introduced for more efficiently solving the following convex problem (Dong et al., 2019; Lake & Tenenbaum, 2010):
${\max_{\Theta}}:F(\Theta)=\log\det(\Theta)-\frac{1}{M}Tr({X^{\top}\Theta X})-\beta{{\|\Theta\|}}_{1},$ (11)
where $\Theta={L}+\frac{1}{\sigma^{2}}I$, ${L}$ denotes the set of valid graph Laplacian matrices, $I$ denotes the identity matrix, and $\sigma^{2}>0$ denotes prior feature variance. It can be shown that the three terms in (11) are corresponding to $\log\det(\Theta)$, $Tr(\Theta S)$ and ${\beta}{{\|\Theta\|}}_{1}$ in (10), respectively. Note that the second term also promotes graph sparsity, so the ${\beta}{{\|\Theta\|}}_{1}$ can be dropped without impacting the final solution. Since $\Theta={L}+\frac{1}{\sigma^{2}}I$ correspond to symmetric and positive definite (PSD) matrices (or M matrices) with non-positive off-diagonal entries, this formulation will lead to the estimation of attractive GMRFs (Dong et al., 2019; Slawski & Hein, 2015). In case $X$ is non-Gaussian, formulation (11) can be understood as Laplacian estimation based on minimizing the Bregman divergence between positive definite matrices induced by the function $\Theta\mapsto-\log\det(\Theta)$ (Slawski & Hein, 2015).
A.4 Proof for Theorem 3.2
In the SAGMAN framework, we aim to construct low-dimensional manifolds for both the input and output of a GNN. For the input manifold, we use the embedding matrix $X=U_{k}\in\mathbb{R}^{|V|\times k}$ (as defined in Definition 3.1), where $|V|$ is the number of nodes in the graph and $k$ is the dimensionality of the embedding space. The goal is to learn a precision matrix $\Theta$ that captures the underlying graph structure reflected in the embeddings.
Maximum Likelihood Estimation (MLE) of the PGM (Precision Matrix).
We start by formulating the MLE of the precision matrix $\Theta$ as a convex optimization problem (Dong et al., 2019):
$\max_{\Theta}\quad F(\Theta)=\log\det(\Theta)-\frac{1}{k}\operatorname{Tr}(X^{\top}\Theta X),$ (12)
where:
  • $\Theta=\mathcal{L}+\frac{1}{\sigma^{2}}I$,
  • $\mathcal{L}$ is the graph Laplacian matrix,
  • $I$ is the identity matrix,
  • $\sigma^{2}>0$ is a prior variance term,
  • $\operatorname{Tr}(\cdot)$ denotes the trace of a matrix.
Expanding the Laplacian Matrix.
The graph Laplacian matrix $\mathcal{L}$ can be decomposed as:
$\mathcal{L}=\sum_{(p,q)\in E}w_{p,q}e_{p,q}e_{p,q}^{\top},$ (13)
where:
  • $E$ is the set of edges in the graph,
  • $w_{p,q}$ is the weight of edge $(p,q)$,
  • $e_{p,q}=e_{p}-e_{q}$, with $e_{p}$ being the standard basis vector corresponding to node $p$.
The edge weights are defined as:
$w_{p,q}=\frac{1}{\|X^{\top}e_{p,q}\|_{2}^{2}}=\frac{1}{\|X_{p}-X_{q}\|_{2}^{2}},$ (14)
where $X_{p}$ and $X_{q}$ are the embeddings of nodes $p$ and $q$, respectively.
Expressing the Objective Function.
Substituting $\Theta$ and $\mathcal{L}$ into the objective function, we can split $F(\Theta)$ into two parts:
$F=F_{1}-\frac{1}{k}F_{2},$ (15)
where:
  1. 1.
    First Term ($F_{1}$):
    $F_{1}=\log\det(\Theta)=\sum_{i=1}^{|V|}\log\left(\lambda_{i}+\frac{1}{\sigma^{2}}\right),$ (16)
    with $\lambda_{i}$ being the $i$-th eigenvalue of the Laplacian $\mathcal{L}$.
  2. 2.
    Second Term ($F_{2}$):
    $F_{2}=\operatorname{Tr}(X^{\top}\Theta X)=\frac{\operatorname{Tr}(X^{\top}X)}{\sigma^{2}}+\sum_{(p,q)\in E}w_{p,q}\|X^{\top}e_{p,q}\|_{2}^{2}.$ (17)
    The term $\frac{\operatorname{Tr}(X^{\top}X)}{\sigma^{2}}$ is constant with respect to $w_{p,q}$ and can be ignored for optimization over $w_{p,q}$.
Computing Partial Derivatives.
To optimize $F$ with respect to the edge weights $w_{p,q}$, we compute the partial derivatives of $F_{1}$ and $F_{2}$:
Derivative of $F_{1}$ with respect to edge weight:
$\frac{\partial F_{1}}{\partial w_{p,q}}=\sum_{i=1}^{|V|}\frac{1}{\lambda_{i}+\frac{1}{\sigma^{2}}}\frac{\partial\lambda_{i}}{\partial w_{p,q}}.$ (18)
Since $\frac{\partial\lambda_{i}}{\partial w_{p,q}}=v_{i}^{\top}\frac{\partial\mathcal{L}}{\partial w_{p,q}}v_{i}$, where $v_{i}$ is the eigenvector corresponding to $\lambda_{i}$, and $\frac{\partial\mathcal{L}}{\partial w_{p,q}}=e_{p,q}e_{p,q}^{\top}$, we have:
$\frac{\partial F_{1}}{\partial w_{p,q}}=\sum_{i=1}^{|V|}\frac{(v_{i}^{\top}e_{p,q})^{2}}{\lambda_{i}+\frac{1}{\sigma^{2}}}.$ (19)
When $\sigma$ approaches infinity, this expression is known as the effective resistance distance between nodes $p$ and $q$:
$d^{\mathrm{eff}}(p,q)=e_{p,q}^{\top}\Theta^{-1}e_{p,q}.$ (20)
So,
$\frac{\partial F_{1}}{\partial w_{p,q}}=d^{\mathrm{eff}}(p,q).$ (21)
Derivative of $F_{2}$ with respect to edge weight:
$\frac{\partial F_{2}}{\partial w_{p,q}}=\|X^{\top}e_{p,q}\|_{2}^{2}=\|X_{p}-X_{q}\|_{2}^{2}.$ (22)
This is the data distance between nodes $p$ and $q$:
$d^{\mathrm{dat}}(p,q)=\|X_{p}-X_{q}\|_{2}^{2}=\frac{1}{w_{p,q}}.$ (23)
Since $w_{p,q}=\frac{1}{d^{\mathrm{dat}}(p,q)}$, we have:
$\frac{\partial F_{2}}{\partial w_{p,q}}=\frac{1}{w_{p,q}}.$ (24)
Optimization Strategy.
The gradient of $F$ with respect to $w_{p,q}$ is:
$\frac{\partial F}{\partial w_{p,q}}=\frac{\partial F_{1}}{\partial w_{p,q}}-\frac{1}{k}\frac{\partial F_{2}}{\partial w_{p,q}}=d^{\mathrm{eff}}(p,q)-\frac{1}{k}\frac{1}{w_{p,q}}.$ (25)
To maximize $F$, we can:
  • Increase $w_{p,q}$ when $d^{\mathrm{eff}}(p,q)>\dfrac{1}{k}\dfrac{1}{w_{p,q}}$.
  • Decrease $w_{p,q}$ when $d^{\mathrm{eff}}(p,q)<\dfrac{1}{k}\dfrac{1}{w_{p,q}}$.
However, since $w_{p,q}\geq 0$, decreasing $w_{p,q}$ effectively means pruning the edge $(p,q)$.
Edge Pruning Strategy.
We aim to prune edges where:
  • Effective Resistance Distance is Small ($d^{\mathrm{eff}}(p,q)$ is small): The nodes are well-connected in terms of the graph structure.
  • Data Distance is Large ($d^{\mathrm{dat}}(p,q)$ is large): The embeddings of the nodes are far apart.
This leads us to consider the distance ratio $\rho_{p,q}$:
$\rho_{p,q}=\frac{d^{\mathrm{eff}}(p,q)}{d^{\mathrm{dat}}(p,q)}=w_{p,q}\,d^{\mathrm{eff}}(p,q).$ (26)
Edges with small $\rho_{p,q}$ are candidates for pruning because they contribute less to maximizing $F$. By pruning such edges, we focus on retaining edges that have a significant impact on the graph’s spectral properties.
Connection to Spectral Graph Sparsification.
The ratio $\rho_{p,q}$ corresponds to the edge sampling probability used in spectral graph sparsification (Spielman & Teng, 2011). Spectral sparsification aims to approximate the original graph with a sparser graph that preserves its spectral (Laplacian) properties.
In spectral sparsification:
  • Edges are sampled with probability proportional to $w_{p,q}d^{\mathrm{eff}}(p,q)$.
  • Edges with higher $\rho_{p,q}$ are more likely to be included in the sparsified graph.
Therefore, our edge pruning strategy—performing spectral sparsification on the initial graph—is equivalent to maximize the objective function in Equation 2. This ensures that the essential structural properties of the graph are maintained while reducing complexity.
A.5 Experimental Setup
See Appendix A.10 for detailed descriptions of all graph datasets used in this work. We employ the most popular backbone GNN models including GCN (Kipf & Welling, 2016), GPRGNN (Chien et al., 2020), GAT (Veličković et al., 2017), APPNP (Gasteiger et al., 2018), ChebNet (Defferrard et al., 2016), and Polynormer (Deng et al., 2024). The recommendation system is based on PinSage (Ying et al., 2018). Perturbations include Gaussian noise evasion attacks and adversarial attacks (DICE (Waniek et al., 2018), Nettack (Zügner et al., 2018), and FGA (Chen et al., 2018)). The input graph-based manifolds are constructed using graph adjacency and node features, while the output graph-based manifold is created using post-softmax vectors. To showcase SAGMAN’s effectiveness in differentiating stable from unstable nodes, we apply SAGMAN to select $1\%$ of the entire dataset as stable nodes, and another $1\%$ as unstable nodes. This decision stems from that only a portion of the dataset significantly impacts model stability (Cheng et al., 2021; Hua et al., 2021; Chang et al., 2017). Additional evaluation results on the entire dataset can be found in Appendix A.12. We quantify output perturbations using cosine similarity and Kullback-Leibler divergence (KLD). Additional insights on cosine similarity, KLD, and accuracy can be found in Appendix A.7. For a large-scale dataset “ogbn-arxiv", due to its higher output dimensionality, we focus exclusively on accuracy comparisons. This decision is informed by the KLD estimator’s $n^{-\frac{1}{d}}$ convergence rate (Roldán & Parrondo, 2012), where $n$ is the number of samples and $d$ is the dimension. In this paper, our spectral embedding method consistently utilizes the smallest $50$ eigenpairs for all experiments, unless explicitly stated otherwise.
A.6 Measure DMD without PGMs
Attack Embedded Matrix Cosine Similarity (Stable / Unstable)
DICE level 1 0.89 / 0.89
Gaussian noise 1.0 0.96 / 0.96
Table 7: The stable (unstable) nodes identified based on traditional graph-based manifolds (Belkin & Niyogi, 2003)
As shown in Table 7, we cannot obtain meaningful DMDs with traditional graph-based manifolds (kNN graphs) adopted in the Laplacian Eigenmaps framework (Belkin & Niyogi, 2003).
A.7 Metrics for Assessing GNN Stability
@@
A.8 GNN Stability Analysis without GDR